OpenStudy (anonymous):

-3(4--2)=? but what is the expanded or condensed form?

OpenStudy (anonymous):

what do you think you would do first?

OpenStudy (anonymous):

I know that I have to find the answer and its -18 I think but how do I rewrite it into an expanded form?

OpenStudy (the_fizicx99):

-3(4-(-2))= -3(4+2) = ? -12 -6 = -18

OpenStudy (anonymous):

www.aaamath.com/g4_31dpx.htm. This site shows what expanded from. Hope its what you're looking for.